Can you please help me? The problem is the square root of X squared to the third power equals 9. Can you show me how to work this out step by step?
Answer by checkley71(8403)

(SQRTX^2)^3=9 SQRT OF X^2=X THEN
X^3=9 CUBE ROOT OF X^3=X THEN
X=CUBEROOT9